• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 15:22:27 on July22,2005

    Entry number 149033


    keyword=cavity signals / gate change

    During the controlled access, I looked at the (new) cavity
    monitor signals on a scope. The BCM output is flat and near
    zero at the few mV level -- this looks ok. But all 4 BPM signals
    are oscillating with a frequency of order 1 Mhz and an amplitude
    that swings +/- 100 mV (actually a bit asymmetric, and different
    for each channel). This does not seem reasonable. I've put in
    an inquiry to the electronics experts, Musson et.al.

    Another point: I reduced the gate on the ADC 1881M in ROC4
    slot 20. It was 110 nsec wide, now it is 50 nsec wide. The
    cavity signals are the only thing going to this ADC. This
    adjustment will improve the dynamic range of the cavities
    in the spectrometer DAQ.